Trying to find the slope and "Y"-Intercept of the lines for the Equation listed below. I also need to graph it, if possible please show this process so that I can apply it to further problems. Regards Grace
Sorry I put the wrong problem down for my initial question. GZ
Y=4/3x+1/2

Answer by rfer(16322)

y=mx+b
m=slope
b=y intercept
------------------
y=4/3x+1/2
slope m=4/3
y int=1/2